A hot tub with a surface area of 28 ft^2 is filled with water to a depth of 29 in. What is the bolume of the water in thr tub, in liters? Express your answer to two significant figures and include the appropriate units.


1
Expert's answer
2021-02-10T01:45:16-0500

28 ft=28*0.3048 m=8.5344 m=85.344 dm

2 ft=2*0.3048 m=6.096 dm

29 in=29*0.0254 m=0.505 dm


"V=85.344*6.096*0.505=262.73 dm\land3(L)"